A piece of wire 48cm long is cut into two pieces. The first is 4x cm long and each piece is bent to form a square.
a) Show that A, the sum of the areas of the two square will be given by A= 2x^2 -24x + 144.
b) At what value of x will the lowest value of A occur
c)Hence or otherwise, find the lowest value of A

Answer by htmentor(1343)

The perimeter of the first square is 4x, so the length of each side is x
The perimeter of the second square is 48 - 4x, so the length of each side is 12 - x
A1 = x^2
A2 = (12-x)^2 = 144 - 24x + x^2
A = A1 + A2 = x^2 + 144 - 24x + x^2 = 2x^2 - 24x + 144